Problem 2-14 Calculating Total Cash Flows [LO4] Volbeat Corp. shows the following information on its 2015 income statement: sales $242,000; costs $153,000; other expenses $7,900, depreciation expense $17,700; interest expense $14,100, taxes $17,255, dividends $11,000. In addition, youre told that the firm issued $5,600 in new equity during 2015 and redeemed $4,100 in outstanding long-term debt. a. What is the 2015 operating cash flow? (Do not round intermediate calculations.) Operating cash flow b. What is the 2015 cash flow to creditors? (Do not round intermediate calculations.) Cash flow to creditors c. What is the 2015 cash flow to stockholders? (Do not round intermediate calculations.) Cash flow to stockholders d. If net fixed assets increased by $22,000 during the year, what was the addition to NWC? (Do not round intermediate calculations.) Addition to NWC
Answer #1

Volbeat Corp

  1. 2015 operating cash flow:

Cash flow from operations = sales – (costs + other expenses + taxes)

= 242,000 – (153,000 + 7,900 + 17,255) = $63,845

  • Depreciation is a non-cash expense, hence is not deducted to arrive at cash flow from operations.
  • Interest expense is a financial charge and hence does not form part of operating cash flow calculations.

  1. 2015 cash flow to creditors:

Cash flow to creditors = interest - net new long term debt

= $14,100 – (-4,100) = $18,200

Since the company redeemed $4,100 outstanding long term debt, the net new long term debt is shown as a negative amount.

  1. 2015 Cash flow to stockholders:

Cash flow to stockholders = dividends – new equity issue

Cash flow to stockholders = $11,000 - $5,600 = $5,400

  1. Determination of addition to NWC when net fixed assets increase by $22,000:

Addition to NWC –

Cash flow to assets = cash flow to creditors + cash flow to stockholders

= $18,200 + $5,400 = $23,600

Cash flow to assets = operating cash flow – net capital spending – change in NWC

Operating cash flow = $63,845

Net capital spending = increase in fixed assets + depreciation

= $22,000 + $17,700 = $39,700

Cash flow to assets = 63,845 – 39,700 – 23,600 = $545